# Hallucination Detox: Sensitive Neuron Dropout (SeND) for Large Language Model Training

### v1

@@ 3.3 Efficient EigenScore Approximation
To address the computational complexity of EigenScore calculations, particularly as LLM hidden layer sizes increase and more generations occur at higher temperatures, we develop an approximation method. This approximation, detailed in Algorithm 1, leverages the properties of Spectral Density or Density of States (DOS) to estimate EigenScore without explicitly constructing the covariance matrix. While this approximation provides a general overview of EigenScore trends, it is important to note that the output scales differ: EigenScore ranges from , whereas the approximation, referred to as Efficient EigenScore (EES), outputs values between . Since the spectrum of the matrix is altered to make EES computable and operates on its own scale, EES can be seen as a standalone metric for hallucination detection.
@@
Our objective is to simplify the integral and approximate its value, avoiding the direct computation of the covariance matrix. This approach is intended to mitigate the computational complexity and associated costs of explicitly handling the covariance matrix. Further utilizing Chebyshev Polynomials, DOS, and KPM (as introduced in Appendix B.2), we can simplify the integral mentioned in Equation 7 to , where term in DOS is approximated using Stochastic Trace Estimation and m’th Chebyshev Polynomial coefficient. Appendices B.3 and B.4 provide the derivation of this equation. Note that the simplified integral is ultimately used to approximate the EigenScore of the matrix which is ultimately equivalent to . Performance of EES approximation is closely correlated with that of the original EigenScore and can be seen to closely track the progress of EigenScore through training of Pythia 1B on the HELM dataset in Figure 7.
3.4 How does Efficient EigenScore approximation scale compared to regular EigenScore?
The efficiency of EES is compared to that of the regular EigenScore calculation with respect to scaling matrix sizes. These tests are imperative to the application of our training protocol on LLMs in Section 4, as with increasing model size, we get ever larger matrix sizes to decompose for the EigenScore calculation. To elucidate these scaling effects, we conduct a grid search over two important parameters: Matrix size (Figure 3) and Moments used for EES calculation (Figure B.5). The difference between EES time in comparison to EigenScore when increasing the number of columns and rows is visualized in Figure 3 using a moments value of 20. It is evident that EES provides a significant computational advantage when increasing the number of columns or rows. Remarkably, at matrix size , EES nearly halves the computation time of regular EigenScore calculation at around 4 seconds whereas EigenScore takes approximately 7 seconds to calculate. We can then deduce that given a good enough approximation, EES provides a significant reduction in computational complexity as model and matrix size increase.
4 SEnsitive Neuron Dropout (SeND)
Building on the findings from Section 3.2, and aiming to reduce variance in the factual uncertainty of LLMs during training, this section introduces SeND, an efficient and transferable framework for training LLMs. SeND integrates the EES method discussed in Section 3.3 to enhance computational efficiency while addressing variance in sensitive neuron behavior. By identifying sensitive neurons, which contribute to the oscillatory behavior of hallucinations during training, SeND deterministically drops these neurons based on a small subset of the training data. This approach ensures an increase in the model’s factual certainty by the end of training as explained in Algorithm 2.
@@ 4.1 SeND Experiment Setup
To evaluate SeND, we use Eleuther AI’s Pythia 1B model, continuing its training on specific datasets rather than restarting pretraining to maintain efficiency. We continue the training of the fully trained model on two datasets: HELM, consisting of Wikipedia text (Su et al., 2024), and MedHALT, a medical dataset emulating real-world entrance exam questions (Pal et al., 2023). Due to the importance of factual accuracy in the medical domain, MedHALT was chosen to assess SeND’s impact on hallucination mitigation. Both datasets were tested in two sizes: 200 and 2,000 points (referred to as 2k). SeND implements the EigenScore reduction technique from Section 3.2 and detects Sensitive Neurons using a 3-epoch window on a specialized hallucination tracking dataset. Sensitive Neurons in the penultimate layer are identified based on their variability across epochs and are deterministically dropped for the subsequent 3 training epochs. This dropout process is repeated at each 3-epoch interval until the training loss converges, effectively mitigating hallucination tendencies and refining the model. While we acknowledge that the number of subsequent epochs for dropout could be adjusted, we limit our approach to 3 for ease of computation.
4.2 Performance of SeND on Pythia 1B
The results of training Pythia 1B on HELM and MedHALT 200 are illustrated in Figure 4. To validate that the EES method accurately approximates the EigenScore metric; we compare the model’s progress during training (up to loss convergence) and assess whether the resulting graphs are similar. These results are detailed in Appendix B.6. Upon confirming that EES provides a reliable approximation of the EigenScore, we proceed to compare the performance of Pythia 1B trained using standard training with SeND on HELM and MedHALT 2k (Figure 4). In the case of training on HELM with the regular protocol, we observe results consistent with previous findings: while the model successfully reduces loss, it fails to optimize for hallucination, as evidenced by the increasing EES metric (green line in Figure 4(a)). Conversely, training with SeND reveals a consistent trajectory toward reducing both EES and loss, as depicted by the blue line.
@@ 4.2 Performance of SeND on Pythia 1B
To assess the effectiveness of SeND in comparison to other state-of-the-art factuality metrics, we employ the FactScore metric from Min et al. (2023), which quantifies the factual accuracy of content generated by large language models (LLMs). The fact-checking is conducted using the HELM dataset on the final trained models (1B SeND and 1B Normal Training). A higher FactScore indicates improved factual precision. When evaluated on 100 data points from the HELM dataset, the 1B SeND model achieves a FactScore of 0.07, whereas the 1B Normal Training model attains 0.05, demonstrating a 40% improvement in factual accuracy, even during test time. This highlights the efficacy of SeND in enhancing the factual certainty of the model. Note that SeND is not a replacement for post-hoc methods like RAG (Gao et al., 2024b), but rather to complement them.
The robustness of this training protocol is essential as we aim for it to be applied across many different fields. In light of this, we present the results of finetuning on MedHALT 2k with and without SeND in Figure 4(b). We observe a similar trend in 4(b) as shown in Figure 4(a), where standard finetuning increases the EES score throughout training, showing that the model is not taking into account hallucinations and factuality during its training. In Figure 4(b), there is an improvement in the trajectory of EES as training continues, showing that our model is in fact able to incorporate factuality as a metric to account for during training of the model. Although the decrease in EES is marginal for MedHALT 2k, we can attribute this to the dataset it is being trained on. The relatively small difference observed between the training protocol behaviours could be due to MedHALT 2k data never being seen before the finetuning phase whereas HELM data has been seen. In this case, it may be beneficial to delay the onset of SeND, as high variability between checkpoints on new training instances is expected.
@@ 5 Conclusion & Future Work
In this paper, we presented a protocol to refine the current training methods of LLMs based on experiments showing oscillatory behaviour with respect to hallucinations throughout training (Figure 1). To do this we used the internal states of LLMs, specifically the penultimate layer activations during inference on a specialized dataset. We present an initial method of reducing hallucinations based on the principles of EigenScore metrics introduced by Chen et al. (2024). We showed empirically that our Sensitive Neuron detection method significantly reduces the EigenScore on inference of LLMs throughout various stages of training (Figure 2). Following the success of the Sensitive Neuron method, we moved on to the application of a hallucination reduction method on training. We show through finetuning that we are able to fix the oscillatory behaviour initially seen throughout training and reduce the EES of finetuned models as shown in Figure 4 by modifying the internal mechanics of training with SEnsitive Neuron Dropout. At test time we achieve a 40% increase in FactScore performance, verifying that SeND provides a substantial improvement to current training protocols.
To further advance our work, we plan to scale SeND to larger datasets and models, as current experiments were limited by compute constraints with larger LLMs. Demonstrating SeND’s effectiveness on larger open-source models like Meta’s LLaMA 3.1 (Dubey et al., 2024) will provide crucial evidence for organizations developing state-of-the-art LLMs to incorporate SeND into their training protocols, ultimately improving model safety. Given that SeND targets variance reduction during training, we anticipate even greater gains on larger LLMs, where the higher inherent variance may amplify the regularization effect and lead to more significant improvements.
@@ Appendix A Additional Experiments
A.1 Drastic Embedding Changes leading to Sensitive Neurons
Looking at internal states of the model allows getting a deeper understanding of the dynamics that could be leading to the oscillatory behaviour seen in Figure 1. To do this, we record the net change (Definition 3.2) between checkpoints of the penultimate layer where one checkpoint would be the correct answer and the next would hallucinate. This net change with respect to various different input texts is plotted in Figure 5. It can be observed that there were specific embedding activations that experienced drastically more change relative to the rest of the embeddings. This is the main source of motivation to further define Sensitive Neurons (Definition 3.3).
@@ B.6 EigenScore and EES training trajectories
To demonstrate that our EigenScore approximation method, EES, is a good metric, we record the progress of Pythia 1B finetuning on the HELM dataset using both EigenScore and EES hallucination performance metrics (Figure 7). Albeit a different scale and window, the trajectories, magnitude and shape of the graphs are nearly identical while EES takes only 4 minutes to calculate and EigenScore takes approximately 8, an astounding 2x increase in compute speed. These results show that our metric closely resembles the target metric while greatly reducing the required computational resources.

### v2

@@ 3.3 Efficient EigenScore Approximation
To address the computational complexity of EigenScore calculations, particularly as LLM hidden layer sizes increase, we develop an approximation method. This approximation, detailed in Algorithm 1, leverages the properties of Spectral Density or Density of States (DOS) to estimate EigenScore without explicitly constructing the covariance matrix. While this approximation provides a general overview of EigenScore trends, it is important to note that the output scales differ: EigenScore ranges from , whereas the approximation, referred to as Efficient EigenScore (EES), outputs values between . Since the spectrum of the matrix is altered to make EES computable and operates on its own scale, EES can be seen as a standalone metric for hallucination detection.
@@
Our objective is to simplify the integral and approximate its value, avoiding the direct computation of the covariance matrix. This approach is intended to mitigate the computational complexity and associated costs of explicitly handling the covariance matrix. Further utilizing Chebyshev Polynomials, DOS, and KPM (as introduced in Appendix B.2), we can simplify the integral mentioned in Equation 7 to , where term in DOS is approximated using Stochastic Trace Estimation and m’th Chebyshev Polynomial coefficient. Appendices B.3 and B.4 provide the derivation of this equation. Note that the simplified integral is ultimately used to approximate the EigenScore of the matrix which is ultimately equivalent to . Performance of EES approximation is closely correlated with that of the original EigenScore which can be seen in Figure 10.
3.4 How does EES scale compared to regular EigenScore?
The efficiency of EES is compared to that of the regular EigenScore calculation with respect to scaling matrix sizes. These tests are imperative to the application of our training protocol on increasing LLM sizes in Section 4 due to larger matrix sizes to decompose for the EigenScore calculation. We conduct a grid search over two important parameters: Matrix size (Figure 3) and Moments used for EES calculation (Figure B.5). The difference between EES time in comparison to EigenScore when increasing the number of columns and rows is visualized in Figure 3 using a moments value of 20. It is evident that EES provides a significant computational advantage when increasing the number of columns or rows. Remarkably, at matrix size , EES nearly halves the computation time of regular EigenScore calculation at around 4 seconds whereas EigenScore takes approximately 7 seconds to calculate. We can then deduce that given a good enough approximation, EES provides a significant reduction in computational complexity as model and matrix size increase.
4 Sensitivity Dropout (SenD)
Building on the findings from Section 3.2, and aiming to reduce variance in the factual uncertainty of LLMs during training, this section introduces SenD, an efficient and transferable framework for training LLMs. SenD integrates the EES method discussed in Section 3.3 to enhance computational efficiency while addressing variance in SEI behavior. By identifying SEIs, which contribute to the oscillatory behavior of hallucinations during training, SenD deterministically drops these embedding indices based on a small subset of the training data. This approach ensures an increase in the model’s factual certainty by the end of training as explained in Algorithm 2.
@@ 4.1 SenD Experiment Setup
To evaluate SenD, we use Pythia 1B model (Biderman et al., 2023), Llama 3.2 1B, and Llama 3.1 8B (Dubey et al., 2024) continuing their training on specific datasets rather than restarting pretraining for efficiency. We continually train the models on the following datasets: HELM, consisting of Wikipedia text (Su et al., 2024), MedHALT, a medical dataset emulating real-world entrance exam questions (Pal et al., 2023), and LegalBench consisting of data for reasoning in LLMs (Guha et al., 2023). Note that HELM and MedHALT are specifically designed for hallucination detection/mitigation in LLMs. SenD is trained on all datasets using 200 data points and and 2,000 points (referred to as 2k) for MedHALT due to the medical domain importance. SenD implements the EigenScore reduction technique from Section 3.2 and detects SEIs using a 3- checkpoint window on a specialized hallucination tracking dataset. The distance between checkpoints and the dropout rate are tunable hyperparameters. Given our ablation study in Appendix C, we opt for and Threshold=3 for the experiments. SEIs in the penultimate layer are identified based on their variability across checkpoints and are deterministically dropped for the subsequent 3 training checkpoints. This is repeated at each 3- checkpoint interval until loss convergence, effectively mitigating hallucination tendencies and oscillations.
4.2 Performance of SenD on Pythia and Llama Models
Pythia and Llama training results are illustrated in Figure 4. To validate that EES accurately approximates the EigenScore metric; we compare the model’s progress during training detailed in Appendix B.6. Upon confirming that, we proceed to compare the performance of Pythia 1B, Llama 3.2 1B, and Llama 3.1 8B trained using training without dropout to that of SenD (Figure 4). As shown in the figure and detailed in Appendix D, across all three models and most domains, training with SenD results in a greater reduction in EES during training. In most cases, the final model trained with SenD achieves a lower EES compared to standard training which increases the EES score, demonstrating its effectiveness in reducing hallucination variance and improving the model’s final certainty. In addition, there seems to be better improvements using SenD in LegalBench settings. A possible explanation for SenD’s superior performance on LegalBench compared to the other two domains which are more general is that early in training, the model sees less domain-specific data, leading to better generalization. However, confirming this hypothesis requires further investigation beyond the scope of this work.
@@ 4.2 Performance of SenD on Pythia and Llama Models
To assess the effectiveness of SenD in comparison to other state-of-the-art factuality metrics, and not solely rely on EES for hallucination detection, , we employ the FactScore metric (Min et al., 2023), which quantifies the factual accuracy of content generated by large language models (LLMs) and HaluEval in Summarization setting (Li et al., 2023). The fact-checking for FactScore is conducted using the HELM dataset on the final trained models (Pythia 1B SenD and 1B Normal Training). A higher FactScore indicates improved factual precision and for HaluEval, higher Exact Match, Accuracy, and Correctness imply better matching with ground truth facts. FactScore is done on 100 and 1000 data points subsequently and table 1 presents the results which imply the better performance of the end model trained with SenD up to 40%. This also depicts the correlation between low EES and high FactScore or HaluEval metrics which are other SOTA metrics for hallucination detection.
| Task | Metric | SenD | Normal |
| HaluEval Summarization (LMEval) | Accuracy | 0.016 | 0.014 |
| Correctness | 0.027 | 0.027 | |
| Exact Match | 0.589 | 0.496 | |
| FactScore (100 points) | Score | 0.07 | 0.05 |
| FactScore (1000 points) | Score | 0.08 | 0.06 |
Since SenD is the first method to focus on Hallucinations during the training of LLMs, there are no baselines or SOTA methods to compare it to. However, one could treat SenD as a post-hoc method and compare it to RAG (Lewis et al., 2021) to get a sense of how different training based methods compare to in-context methods. When tested on 1000 data points, SenD-continually-trained Pythia achieves a FactScore of 0.07 as in Table 1, while the base Pythia with RAG scores 0.25, highlighting RAG’s effectiveness in reducing hallucinations by providing context. However, to make the comparison fair, applying RAG to a SenD-trained model achieves a higher FactScore of 0.28, outperforming RAG on a normally trained model by 12%. This indicates that even though SenD is not meant to outperform post-hoc methods, SeND combined with RAG enhances the end model’s factual performance compared to RAG on a normally trained model.
@@ 5 Conclusion & Future Work
In this paper, we presented a protocol to refine the current training methods of LLMs based on experiments showing oscillatory behaviour with respect to hallucinations throughout training (Figure 1). To do this we used the internal states of LLMs, specifically the penultimate layer activations during inference on a specialized dataset. We present an initial method of reducing hallucinations based on the principles of EigenScore metrics introduced by Chen et al. (2024). We showed empirically that our SEI detection method significantly reduces the EigenScore on inference of LLMs throughout various stages of training (Figure 2). Following the success of the SEI method, we moved on to the application of a hallucination reduction method on training on Pythia and Llama models and various domains. We show through training with SenD that we are able to fix the oscillatory behaviour initially seen throughout training and reduce the EES of finetuned models as shown in Figure 4 by modifying the internal mechanics of training with Sensitivity Dropout. At test time we achieve a 40% increase in FactScore performance and improvement of other SOTA hallucination detection metrics, verifying that SenD provides a substantial improvement to current training protocols both during and after training.
Note that although SenD has only been applied to finetuning training in this paper, the training framework is applicable to all stages of training. We encourage future work to implement SenD on larger training sets, such as pretraining, to see how SenD performs in these environments. To further advance our work, we plan to scale SenD to larger datasets and models, as current experiments were limited by compute constraints with larger LLMs. Demonstrating SenD’s effectiveness on larger open-source models like Meta’s LLaMA 3.2 405B (Dubey et al., 2024) will provide crucial evidence for organizations developing state-of-the-art LLMs to incorporate SenD into their training protocols, ultimately improving model safety. Given that SenD targets variance reduction during training, we anticipate even greater gains on larger LLMs, where the higher inherent variance may amplify the regularization effect and lead to more significant improvements.
@@ Appendix A Additional Experiments
A.1 Drastic Embedding Changes leading to Sensitive Embedding Indices
Looking at internal states of the model allows getting a deeper understanding of the dynamics that could be leading to the oscillatory behaviour seen in Figure 1. To do this, we record the net change (Definition 3.2) between checkpoints of the penultimate layer where one checkpoint would be the correct answer and the next would hallucinate. This net change with respect to various different input texts is plotted in Figure 5. It can be observed that there were specific embedding activations that experienced drastically more change relative to the rest of the embeddings. This is the main source of motivation to further define SEIs (Definition 3.3).
A.2 Hallucination Oscillations Across Model Sizes
Figures 6, 7, and 8 show our study of hallucination oscillations during the training of Pythia models. An overall observation across the plots is that as opposed to our intuitive expectation which is a linear decrease of the hallucination detection metric when the model scales linearly, neither the oscillations during the training of the model decrease, nor the end model reaches its optimal state in terms of the hallucination metric.
@@ B.6 EigenScore and EES training trajectories
To demonstrate the high correlation between EigenScore and EES, we record the progress of Pythia 1B finetuning on the HELM dataset using both EigenScore and EES hallucination metrics (Figure 10). Albeit a different scale and window, the trajectories, magnitude and shape of the graphs are nearly identical while EES takes only 4 minutes to calculate and EigenScore takes approximately 8, an astounding 2x increase in compute speed. These results show that our metric closely resembles the target metric while greatly reducing the required computational resources.
Appendix C Ablation study on K and Step Thresholding for SenD
Figure 11 shows the ablation study done on and Figure 12 illustrates the ablations study done on the Step Threshold for SenD experiments. As depicted, and Threshold = 3 are chosen for our experiments except for Llama 3.1 8B model which due to its larger size requires more embedding indices to be dropped, hence adapting to .
Appendix D Additional Pythia 1B, Llama 3.2 1B, and Llama 3.1 8B Training with and without SenD
Here, we present additional experimental results of training Pythia and LLaMA on multiple domains. Figure 14(b) supplements the results discussed in Section 4 by illustrating the training procedures on the Medical domain. In the Pythia 1B setting, the EES achieved with training using SeND remains consistently lower than that of normal training and exhibits fewer oscillations throughout the training process. In the LLaMA 3.1 8B setting, while both approaches show an increase in the EES metric during training, the final model trained with SeND achieves a lower EES, indicating a reduced likelihood of hallucinations in this domain. Figure 14 visualizes the training results for the LLaMA 1B model across all domains. Similar trends to the other experimental settings are observed. In the HELM and LegalBench settings, the EES of the model trained with SeND shows a significant reduction by the end of the training. However, in the MedHalt setting, this reduction is less pronounced by the conclusion of training. Figure 15 depicts the standard deviations of 5 runs experimented on the Pythia 1B model. As shown, the error shading area is not too wide hence proving the reliability of the results.